# Effectively Leveraging Momentum Terms in Stochastic Line Search Frameworks for Fast Optimization of Finite-Sum Problems

### v1

@@ 4 The Algorithmic Framework
The conceptual distinction between heavy-ball and CG is subtle and not unambiguously recognized by the optimization community. There is a substantial equivalence in the convex quadratic case, if heavy-ball parameters are chosen optimally [28]. In the general nonlinear case, CG methods define the search direction by means of a predetermined formula for $\beta_{k}$ and then carry out a Wolfe line search; on the other hand, heavy-ball selects $\alpha_{k}$ and $\beta_{k}$ altogether by some rule. If an Armijo-type line search is employed after choosing $\alpha_{k}$ and $\beta_{k}$ according to some rule, as done, e.g., in [20, 24], we are arguably somewhere in a middle ground. We propose a framework that belongs to this middle ground in the stochastic setting.
@@
A conjugate-gradient method for finite-sum problem with convergence guarantees has been proposed recently in the literature [34]. The algorithm is substantially a direct descendant of classical CG methods in nonlinear optimization and is shown to possess convergence properties in expectation. However, it presents two major differences w.r.t. what we present here, which actually make it very unappealing in deep learning scenarios: it requires a Wolfe-type line search, significantly increasing the per-iteration cost, and it employs variance reduction techniques, which have a massive impact on the memory footprint of the method.
@@
Several formulae have been proposed that can be used to compute $\beta_{k}$, that collapse to the same rule when $f$ is quadratic [28]. Given $y_{k}=\nabla f(x^{k})-\nabla f(x^{k-1})$, some well-known formulations for computing $\beta_{k}$ (cfr. [22, Sec. 12.5]) are:
$\beta_{k}^{HS}=\frac{\nabla f(x^{k})^{T}y_{k}}{(x^{k}-x^{k-1})^{T}y_{k}},$ (Hestenes-Stiefel)
$\beta_{k}^{FR}=\frac{||\nabla f(x^{k})||^{2}}{||\nabla f(x^{k-1})||^{2}},$ (Fletcher-Reeves)
$\beta_{k}^{PPR}=\frac{\nabla f(x^{k})^{T}y_{k}}{||\nabla f(x^{k-1})||^{2}}.$ (Polyak-Polak-Ribiére)
We shall note that, with the above rules, we exploit the information available during iteration $k-1$ to define the coefficient $\beta_{k}$ for iteration $k$. Thus, in a stochastic scenario, rules of this kind would only be reasonable if $f_{k}$ did not change too much from an iteration to the other. This condition somewhat becomes true under mini-batch persistency.
Indeed, let ${B}_{k}\cap{B}_{k+1}=\mathcal{S}_{k}\neq\emptyset$ and $f_{\mathcal{S}_{k}}=\frac{1}{|\mathcal{S}_{k}|}\sum_{i\in{\mathcal{S}_{k}}}f_{i}(x)$. We can use the values of $f_{\mathcal{S}_{k}}(x^{k})$, $f_{\mathcal{S}_{k}}(x^{k+1})$, $\nabla f_{\mathcal{S}_{k}}(x^{k})$ and $\nabla f_{\mathcal{S}_{k}}(x^{k+1})$, that are available during iteration $k$, to compute the value $\beta_{k+1}$ according to one of the above formulae. The momentum term at iteration $k+1$ will thus be suitably selected for the function $f_{\mathcal{S}_{k}}$, which actually constitutes a substantial portion of the function $f_{k+1}$ to be optimized at iteration $k+1$.
The advantage of this strategy is twofold: not only the estimated $\beta_{k}$ will be meaningful when the mini-batch changes, but we also do not need to make additional evaluations, since the four quantities involved in the CG-type formulae would be computed anyway throughout the algorithm.
It is clear that the proposed strategy, although reasonable, has a heuristic nature, since $\beta_{k}$ is estimated without looking at a portion (say a half) of currently considered data. We thus need to ensure that the obtained direction is actually of descent for $f_{k}$, before starting a line search procedure along the direction. Clearly, if $x^{k}-x^{k-1}$ is a direction of descent for $f_{k}$, i.e., $(x^{k}-x^{k-1})^{T}\nabla f_{k}(x^{k})<0$, the descent property will hold for $d_{k}$ for any choice of $\beta_{k}\geq 0$. From what we observed in the numerical results in Section 3.2, this will often be the case in practice. Yet, in the unfortunate case the direction $d$ obtained by the CG-type formula does not satisfy the above conditions, we can remedy by resorting to any of the following recovery strategies:
  • switch to the negative stochastic gradient $-g_{k}(x^{k})=-\nabla f_{k}(x^{k})$ as a search direction;
@@
  • employ the damping strategy from [20];
  • compute the direction according to the strategy proposed in [24], i.e.,
@@
    where the sequence of matrices $\{H_{k}\}\subseteq\mathbb{R}^{2\times 2}$ satisfies suitable uniformly boundedness assumptions.
The complete pseudocode of the proposed method is reported in Algorithm 1. A few aspects are worth to be remarked about the algorithm:
@@
    In practical implementations, the algorithm will often be designed to work in a random-reshuffling fashion [33]. However, for the sake of simplicity and generality, we describe it in a “pure” stochastic setup.
@@
    We do not specify the particular way of choosing the tentative step size $\alpha_{0}^{k}$ at line 14. There are various options that might work both theoretically and computationally. Of course, a fixed value could be used for all iterations; however, we might prefer to define it adaptively. For this latter approach, a heuristic method, such as the one proposed by Vaswani et al. [17], could be considered, or we could use the generalized Stochastic Polyak step size introduced by Wang et al. [19].
@@
    Finally, we shall underline that Algorithm 1 is shown to be using a monotone Armijo-type line search at step 15. This is done to ease the description of the method, but nothing prohibits to employ a nonmonote line search. Similarly, we explicitly used the Fletcher-Reeves type formula at step 17, but any other CG rule could be employed in principle.
@@
1:Input: $x^{0}\in\mathbb{R}^{n}$, $f_{1},\dots,f_{N}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $\delta\in(0,1)$, $\gamma\in(0,1)$, $M\geq 2$, $c_{1},c_{2}>0$.
2:Let $\mathcal{S}_{0}=\emptyset$,
3:$\beta_{1}=0$, $d_{0}=0$
@@
5:     Randomly draw a subset $\mathcal{S}_{k}\subset\{1,\ldots,N\}$ of size $M/2$
6:     Set $B_{k}=\mathcal{S}_{k-1}\cup\mathcal{S}_{k}$
@@
15:     Set $\alpha_{k}=\max_{j=0,1,\ldots}\{\alpha_{0}^{k}\delta^{j}\mid f_{k}(x_{k}+\alpha_{0}^{k}\delta^{j}d_{k})\leq f_{k}(x_{k})+\gamma\alpha_{0}^{k}\delta^{j}\nabla f_{k}(x_{k})^{T}d_{k}\}$

### v3

@@ 4 Data-persistent Conjugate Gradient Rules
The conceptual distinction between heavy-ball and CG is subtle and not unambiguously recognized by the optimization community. There is a substantial equivalence in the convex quadratic case, if heavy-ball parameters are chosen optimally [polyak1987introduction]. In the general nonlinear case, CG methods define the search direction by means of a predetermined formula for $\beta_{k}$ and then carry out a Wolfe line search; on the other hand, heavy-ball selects $\alpha_{k}$ and $\beta_{k}$ altogether by some rule. If an Armijo-type line search is employed after choosing $\alpha_{k}$ and $\beta_{k}$ according to some rule, as done, e.g., in [fan2023, lapucci2024globally], we are arguably somewhere in a middle ground. We propose a framework that belongs to this middle ground in the stochastic setting.
@@
In the recent literature, two conjugate-gradient methods for finite-sum problem with convergence guarantees have been proposed [jin2019stochastic, kou2023mini]. The two algorithms are substantially direct descendants of classical CG methods in nonlinear optimization, both possessing convergence properties in expectation. However, they present two major differences w.r.t. what we present here, which actually make their employment very unappealing in deep learning scenarios: they require a Wolfe-type line search, significantly increasing the per-iteration cost, and they employ variance reduction techniques, as the algorithm presented in [jin2019stochastic] requires periodic full batch evaluations of the gradient, and the one presented in [kou2023mini] stores per sample gradients, which have a massive impact on the memory footprint of the method.
@@
Several formulae have been proposed that can be used to compute $\beta_{k}$, that collapse to the same rule when $f$ is quadratic [polyak1987introduction]. Given $y_{k}=\nabla f(x^{k})-\nabla f(x^{k-1})$, some well-known formulations for computing $\beta_{k}$ (cfr. [grippo2023introduction, Sec. 12.5]) are:
$\beta_{k}^{HS}=\frac{\nabla f(x^{k})^{T}y_{k}}{(x^{k}-x^{k-1})^{T}y_{k}},$ (Hestenes-Stiefel) (11)
[…]
We shall note that, with the above rules, we exploit the information available during iteration k1k-1 to define the coefficient $\beta_{k}$ for iteration $k$. Thus, in a stochastic scenario, rules of this kind would only be reasonable if $f_{k}$ did not change too much from an iteration to the other. This condition somewhat becomes true under mini-batch persistency.
Indeed, let ${B}_{k}\cap{B}_{k+1}=\mathcal{R}_{k}\neq\emptyset$ and
$f_{\mathcal{R}_{k}}=\frac{1}{|\mathcal{R}_{k}|}\sum_{i\in{\mathcal{R}_{k}}}f_{i}(x).$
We can use the values of $\nabla f_{\mathcal{R}_{k}}(x^{k})$ and $\nabla f_{\mathcal{R}_{k}}(x^{k+1})$, that are available during iteration $k$, to compute the value $\beta_{k+1}$ according to one of the above formulae. For instance, Fletcher-Reeves update becomes
$\beta_{k+1}=\frac{\|\nabla f_{\mathcal{R}_{k}}(x^{k+1})\|^{2}}{\|\nabla f_{\mathcal{R}_{k}}(x^{k})\|^{2}}.$
The momentum term at iteration $k+1$ will thus be suitably selected for the function $f_{\mathcal{R}_{k}}$, which actually constitutes a substantial portion of the function $f_{k+1}$ to be optimized at iteration $k+1$.
The advantage of this strategy is twofold: not only the estimated $\beta_{k}$ will be meaningful when the mini-batch changes, but we also do not need to make additional evaluations, since the four quantities involved in the CG-type formulae would be computed anyway throughout the algorithm. Note that using $\mathcal{R}_{k}$ to compute $\beta_{k+1}$ substantially differs from the strategy used in [jin2019stochastic, kou2023mini], as this allows to compute the CG update rule exactly for the function $f_{\mathcal{R}_{k}}$.
5 The Algorithmic Framework
It is clear that the strategy presented in Section 4, although reasonable, has a heuristic nature, since $\beta_{k}$ is estimated without looking at a portion (say a half) of currently considered data. To obtain a sound algorithm, we must ensure that the obtained direction is actually of descent for $f_{k}$, before starting a line search procedure along the direction: if that were not the case, the backtracking algorithm could get stuck in an infinite loop.
Clearly, if $x^{k}-x^{k-1}$ is a direction of descent for $f_{k}$, i.e., $(x^{k}-x^{k-1})^{T}\nabla f_{k}(x^{k})<0$, the descent property will hold for $d_{k}$ for any choice of $\beta_{k}\geq 0$. From what we observed in the numerical results in Section 3.2, this will often be the case in practice. Yet, in the unfortunate case the direction $d$ obtained by the CG-type formula does not satisfy the above conditions, we shall resort to any of the following recovery strategies:
  • switch to the negative stochastic gradient $g_{k}(x^{k})=-\nabla f_{k}(x^{k})$ as a search direction;
@@ 5 The Algorithmic Framework
  • employ the damping strategy from [fan2023];
  • compute the direction according to the strategy proposed in [lapucci2024globally], i.e.,
@@
    where the sequence of matrices $\{H_{k}\}\subseteq\mathbb{R}^{2\times 2}$ satisfies suitable uniform boundedness assumptions.
The complete pseudocode of the resulting proposed method is reported in Algorithm 1. A few aspects are worth to be remarked about the algorithm:
@@
    In practical implementations, the algorithm will often be designed to work in a random-reshuffling fashion [gurbuzbalaban2021random]. However, for the sake of simplicity and generality, we describe it in a “pure” stochastic setup.
@@
    The set $\mathcal{R}_{k}$ identifies the training samples that are reused in consecutive mini-batches, i.e. $B_{k}\cap B_{k+1}=\mathcal{R}_{k}$. In practice, at line 17 of the algorithm, we can build $\mathcal{R}_{k}$ selecting from $B_{k}$ the indices that were drawn in the most recent iterations at step 5, so that during an epoch we avoid using the same training samples for many iterations in a row. For example, when using a $50\%$ overlap, we will have $\mathcal{R}_{k}=\mathcal{S}_{k}$ so as to use each training sample twice per epoch. As we highlight in Section 5.1, using a deterministic rule at line 17 is crucial to characterize the bias of the gradient estimate $g_{k}(x^{k})$.
@@
    We do not specify the particular way of choosing the tentative step size $\alpha_{0}^{k}$ at line 14. There are various options that might work both theoretically and computationally. Of course, a fixed value could be used for all iterations; however, we might prefer to define it adaptively. For this latter approach, a heuristic method, such as the one proposed by Vaswani et al. [vaswani2019], could be considered, or we could use the generalized Stochastic Polyak step size introduced by Wang et al. [wang2023].
  5. 5.
    Finally, we shall underline that Algorithm 1 is shown to be using a monotone Armijo-type line search at step 15. This is done to ease the description of the method, but nothing prohibits to employ a nonmonotone line search.
@@
1:Input: $x^{0}\in\mathbb{R}^{n}$, $f_{1},\dots,f_{N}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $\gamma\in(0,1)$, $\delta\in(0,1)$.
2:Let $\mathcal{R}_{0}=\emptyset$,
[…]
@@
5:  Randomly draw a subset $\mathcal{S}_{k}\subset\{1,\ldots,N\}\setminus\mathcal{R}_{k-1}$
6:  Set $B_{k}=\mathcal{R}_{k-1}\cup\mathcal{S}_{k}$
@@
15:  Set $\alpha_{k}=\max_{j=0,1,\ldots}\{\alpha_{0}^{k}\delta^{j}\mid f_{k}(x_{k}+\alpha_{0}^{k}\delta^{j}d_{k})\leq f_{k}(x_{k})+\gamma\alpha_{0}^{k}\delta^{j}g_{k}(x^{k})^{T}d_{k}\}$